Silverton Co. is comparing two different capital structures. Plan I would result in 8,000 shares of stock and $456,000 in debt. Plan II would result in 13,700 shares of stock and $239,400 in debt. The interest rate on the debt is 11 percent.

a. Ignoring taxes, compare both of these plans to an all-equity plan assuming that EBIT will be $54,800. The all-equity plan would result in 20,000 shares of stock outstanding. Compute the EPS for each plan. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

EPS
Plan I $
Plan II $
All-equity plan $


b. In part (a), what is the break-even level of EBIT for Plan I as compared to that for an all-equity plan? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

EBIT            $

In part (a), what is the break-even level of EBIT for Plan II as compared to that for an all-equity plan? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

EBIT            $

c. Ignoring taxes, at what level of EBIT will EPS be identical for Plans I and II? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

EBIT            $

d. Assume the corporate tax rate is 32 percent.

Compute the EPS for each plan. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

EPS
Plan I $
Plan II $
All-equity plan $


What is the break-even level of EBIT for Plan I as compared to that for an all-equity plan? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

EBIT            $

What is the break-even level of EBIT for Plan II as compared to that for an all-equity plan? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

EBIT            $

At what level of EBIT will EPS be identical for Plans I and II? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

EBIT            $
Answer #1

1.
=(54800-11%*456000)*(1-0%)/8000=0.58

2.
=(54800-11%*239400)*(1-0%)/13700=2.0778102189781

3.
=(54800-11%*0)*(1-0%)/20000=2.74

4.
=(11%*456000-11%*0*8000/20000)/(1-8000/20000)=83600

5.
=(11%*239400-11%*0*13700/20000)/(1-13700/20000)=83600

6.
=(11%*456000-11%*239400*8000/13700)/(1-8000/13700)=83600

7.
=(54800-11%*456000)*(1-32%)/8000=0.3944

8.
=(54800-11%*239400)*(1-32%)/13700=1.41291094890511

9.
=(54800-11%*0)*(1-32%)/20000=1.8632

10.
=(11%*456000-11%*0*8000/20000)/(1-8000/20000)=83600

11.
=(11%*239400-11%*0*13700/20000)/(1-13700/20000)=83600

12.
=(11%*456000-11%*239400*8000/13700)/(1-8000/13700)=83600